40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 13010     Entry time: Tue May 23 22:58:23 2017     In reply to: 13003     Reply to this: 13015
Author: gautam 
Type: Update 
Category: General 
Subject: De-Whitening board noises 

Summary:

I wanted to match a noise model to noise measurement for the coil-driver de-whitening boards. The main objectives were:

  1. Make sure the various poles/zeros of the Bi-Quad stages and the output stage were as expected from the schematics
  2. Figure out which components are dominating the noise contribution, so that these can be prioritized while swapping out the existing thick-film resistors on the board for lower noise thin-film ones
  3. Compare the noise performance of the existing configuration, which uses an LT1128 op-amp (max output current ~20mA) to drive the input of the coil-driver board, with that when we use a TLE2027 (max output current ~50mA) instead. This last change is motivated by the fact that an earlier noise-simulation suggested that the Johnson noise of the 1kohm input resistor on the coil driver board was one of the major noise contributors in the de-whitening board + coil driver board signal chain. Since the TLE2027 can drive an output current of up to 300mA, we could reduce the input impedance of the coil-driver board to mitigate this noise source to some extent. 

Measurement:

  • The back-plane pin controlling the MAX333A that determines whether de-whitening is engaged or not (P1A) was pulled to ground (by means of one of the new extender boards given to us by Ben Abbott). So two de-whitening stages were engaged for subsequent tests.
  • I first measured the transfer function of the signal path with whitening engaged, and then fit my LISO model to the measurement to tweak the values of the various components. This fitted file is what I used for subsequent noise analysis. 
  • ​For the noise measurement, I shorted the input of the de-whitening board (10-pin IDE connector) directly to ground.
  • I then measured the voltage noise at the front-panel SMA connector with the SR785
  • The measurements were only done for 1 channel (CH1, which is the UL coil) for 4 de-whitening boards (2 ITMs, BS, and SRM). The 2 ITM boards are basically identical, and the BS and SRM boards are similar. Here, only results for the board labelled "ITMX" are presented.
  • For this board, I also measured the output voltage noise when the LT1128 was replaced with a TLE2027 (SOIC package, soldered onto a SOIC-to-DIP adaptor). Steve has found (ordered?) some DIP variants of this IC, so we can compare its noise performance when we get it.

Results:

  • Attachment #1 shows the modeled and measured noises, which are in fairly good agreement.
  • The transfer function measurement/fitting (not attached) also suggests that the poles/zeros in the signal path are where we expect as per the schematic. I had already verified the various resistances, but now we can be confident that the capacitance values on the schematic are also correct. 
  • The LT1128 and TLE2027 show pretty much identical noise performance.
  • The SR785 noise floor was low enough to allow this measurement without any pre-amp in between. 
  • I have identified 3 resistors from the LISO model that dominate the noise (all 3 are in the Bi-Quad stages), which should be the first to be replaced. 
  • There are some pretty large 60 Hz harmonics visible. I thought I was careful enough avoiding any ground loops in the measurement, and I have gotten some more tips from Koji about how to better set up the measurement. This was a real problem when trying to characterize the Coil Driver noise.

Next steps:

  • I have data from the other 3 boards I pulled out, to be updated shortly.
  • The last piece (?) in this puzzle is the coil driver noise - this needs to be modeled and measured.
  • Once the coil driver board has been characterized, we need to decide what changes to make to these boards. Some things that come to mind at the moment:
    • Replace critical resistors (from noise-performance point of view) with low noise thin film ones.
    • Remove the "fast analog" path on the coil driver boards - these have potentiometers in series with the coil, which we should remove since we are not using this path anyways.
    • Remove all AD797s from both de-whitening and coil driver boards - these are mostly employed as monitor points that go to the backplane connector, which we don't use, and so can be removed.
    • Increase the series resistor at the output of the coil driver (currently, these are either 100ohm or 400ohm depending on the optic/channel). I need to double check the limits on the various LSC servos to make sure we can live with the reduced range we will have if we up these resistances to 1 kohm (which serves to reduce the current noise to the coils, which is ultimately what matters).
Attachment 1: ITMX_deWhite_ch1_noise.pdf  262 kB  | Hide | Hide all
ITMX_deWhite_ch1_noise.pdf
ELOG V3.1.3-